OpenStudy (anonymous):

(HELP PLEASE, kind of have trouble towards the end maybe) Jeff and kirk can build a 75 -ft. retaining wall together in 4 hours. because Jeff has more experience, he could build the wall by himself in 1 hour quicker than kirk. How long would it take kirk (to the nearest minute) to build the wall by himself? ( I know kirk takes t+1 (1/t+1) jeff would be 1/t and the eqution would be 4/t + 4/t+1=1 but after this I seem to get lost and get the wrong answer)

OpenStudy (anonymous):

sorry i can't help u

OpenStudy (misty1212):

HI!! do you know what the right answer is ?

OpenStudy (anonymous):

no I don't I tried what I had gotten and it was wrong I only have about two or three chances left.

OpenStudy (misty1212):

if that is really the equation (and it looks like it to me) then we can solve it

OpenStudy (anonymous):

yeah pretty sure it is

OpenStudy (misty1212):

\[\frac{4(2x+1)}{x(x+1)}=1\] or \[4(2x+1)=x^2+x\] solve that quadratic equation, you need the formula

OpenStudy (anonymous):

I got 7.53 and -0.53, this is were I most likely got stuck how to I turn these to into minutes?